A and B are two events. Which of the following holds?

A) P(B|A) is greater than or equal to P(B)

B.) P(A)+P(B) is less than or equal to 1

C.) P(A|B) is greater than or equal to P(A)

D.) none of the answers.
Answer #1

D) None of the answers

Consider the probabilities, P(A) = 0.6, P(B) = 0.7

and P(A & B) = 0.30

P(B | A) = P(A & B)/P(B) = 0.30/0.60 = 0.50 < P(B)

Therefore, statement (A) is wrong.

P(A) + P(B) = 0.6 + 0.7 = 1.3

Therefore, statement (B) is wrong.

P(A | B) = 0.30/0.70 = 0.43 < P(A)

Therefore,statement (C) is wrong.